OpenStudy (anonymous):

Evalute 6n2 − 14n − 36 given n(n − 6) = 16. Smaller Value Larger Value

OpenStudy (anonymous):

Jimmy can you show me how you did the last one? Thanks!!

OpenStudy (anonymous):

same as the last time... you do n^2-6n-16=0 then factorise (x-8)(x+2)=0 then x=8 or -2

OpenStudy (anonymous):

Then you substitute those values into the first equation

OpenStudy (anonymous):

HUH?

OpenStudy (anonymous):

Oh sorry

OpenStudy (anonymous):

um, when you substitute, the answers are

OpenStudy (anonymous):

I can't get the larger number but the smaller number is 16 and I don't know how I got that even

OpenStudy (anonymous):

-52 and and -32

OpenStudy (anonymous):

Oh... well then... am not sure

OpenStudy (anonymous):

-8 and 2 are definitely the substitution factors, however.

OpenStudy (anonymous):

I mean, 8 and -2

OpenStudy (lalaly):

order is right 6n2 − 14n − 36 given n(n − 6) = 16. first thing find the values of n from the second equation n(n-6)=16 n^2-6n-16=0 (n-8)(n+2)=0 n=8, -2 then u substitute n=8 in the first equation 6(8)^2-14(8)-36 and evaluate it then substitute n=-2 in the same equation and evaluate it

OpenStudy (anonymous):

SO, yes, the answers should be -52 and -32... double checked that answer

OpenStudy (anonymous):

is it to the power of 2, or times it by 2?

OpenStudy (lalaly):

power i guess

OpenStudy (anonymous):

Hm. Ok, then that changes the final answers...

OpenStudy (lalaly):

i think whoever askd the question should try to find them

OpenStudy (anonymous):

Ok, well, that's the basics of it, good luck.. I hope you find the final answers! Do you have the answer booklet?